# Non-singular and probability measure-preserving actions of infinite permutation groups

### v1

@@ 3.2. Induced actions
The action does not depend on the choice of the section $s$ (up to conjugacy). Note that if $\sigma$ preserves the measure $\mu$, then $\operatorname{Ind}_{V}^{G}(\sigma)$ preserves the measure $\xi\otimes\mu$. Also, if $\sigma$ is ergodic, so is $\operatorname{Ind}_{V}^{G}(\sigma)$.
@@ 3.3. A classification theorem for non-singular actions
First, we show that there is a non-null $G$-invariant set $E\in\mathrm{MALG}(\mu)$ such that the action of $G$ on $E$ is isomorphic to $\operatorname{Ind}_{V}^{G}(\alpha)$ as above. Let $\mathcal{H}=L^{2}(\Omega,\mu)$ and consider the Koopman representation $\kappa$ of $G$ on $\mathcal{H}$ given by (3.1). Let $(\mathcal{A},\leq)$ and the subspaces $\mathcal{H}_{a}$ for $a\in\mathcal{A}$ be defined as in Subsection 2.2. By Subsection 2.2, $\bigcup_{a\in\mathcal{A}}\mathcal{H}_{a}$ is dense in $\mathcal{H}$, so there exists $a$ with $\mathcal{H}_{a}\neq 0$. By Subsection 2.1, there is a $<$-minimal element $a\in\mathcal{A}$ with $\mathcal{H}_{a}\neq 0$.
@@
Denote the probability measure-preserving action of $G_{a}$ on $(E,\nu_{a})$ by $\alpha$. We will see that the actions $\operatorname{Ind}_{G_{a}}^{G}(\alpha)$ and $G\curvearrowright\bigsqcup_{t\in T}t\cdot E$ are isomorphic. In order to make the isomorphism more transparent, we choose the section $s\colon G/G_{a}\to G$ appearing in the definition of the cocycle (missing) 3.2 to have image $T$, i.e,
@@
To conclude with the proof of the theorem, we apply Zorn’s lemma. Let $\mathcal{D}$ be the collection of all elements of $\mathrm{MALG}(\mu)$ which are $G$-invariant and such that the restriction of the action of $G$ to them is of the form (3.3), and order it by inclusion. If $\mathcal{C}\subseteq\mathcal{D}$ is a chain, we can always find a cofinal set $\mathcal{C}^{\prime}\subseteq\mathcal{C}$, which is countable, because $(\mathcal{C},\subseteq)$ is isomorphic to ($(\{\mu(C):C\in\mathcal{C}\},\leq)\subseteq({\mathbf{R}},\leq)$. It is clear that $\bigcup\mathcal{C}^{\prime}$ is an upper bound for $\mathcal{C}$. Let $D$ be a maximal element of $\mathcal{D}$. If $D\neq\Omega$, we can apply what we already proved to the action of $G$ on $\Omega\setminus D$ and contradict its maximality. ∎
@@ 4. A generalization of de Finetti’s theorem
Let $E\in\mathcal{J}$. It follows from (missing) 4.1 that there exists a finite set $A\subseteq M$ and a $\mathcal{F}_{A}$-measurable event $E^{\prime}$ with $\mu(E\triangle E^{\prime})<\epsilon$. If $B\subseteq M$ is any finite set, by Section 4, there is $g\in G$ with $g\cdot A\cap B=\emptyset$. By invariance of $E$, this implies that for any $B$, there is an $\mathcal{F}_{M\setminus B}$-measurable event $E^{\prime\prime}$ (namely, $E^{\prime\prime}=g\cdot E^{\prime}$) with $\mu(E\triangle E^{\prime\prime})<\epsilon$. Taking limits as $B$ exhausts $M$ and $\epsilon\to 0$, we obtain that $E\in\mathcal{T}$. ∎
@@
[…]
@@
For every $g\in G_{a_{i}}$, we have that $g\cdot\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}=\operatorname{\mathbf{E}}_{g\cdot\mathcal{T}}g\cdot\xi_{i}=\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}$, so $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}\in\mathcal{F}_{a_{i}}$. In particular, $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}$ is $\mathcal{G}$-measurable, so also $\mathcal{T}\cap\mathcal{G}$-measurable. As $\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}\operatorname{\mathbf{E}}_{\mathcal{T}}=\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}$, this implies that $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}=\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}\xi_{1}\cdots\xi_{n}$. Linear combinations of variables of the form $\xi_{1}\cdots\xi_{n}$ generate a dense subspace of $L^{2}(\mathcal{G})$, so $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}$ generate a dense subspace of $L^{2}(\mathcal{T}\cap\mathcal{G})$. However, by (missing) 4.2, each of them belongs to $\bigvee_{a}\mathcal{T}\cap\mathcal{F}_{a}$ and we are done. ∎
@@
Suppose now that missing (i) holds and let the tuple $\bar{b}$ enumerate the set $\bigcup_{n}B_{n}$. By uniform non-algebraicity, there exist $\bar{b}^{\prime}$ and $a^{\prime}\neq a^{\prime\prime}$ with $\operatorname{tp}(\bar{b}a)=\operatorname{tp}(\bar{b}^{\prime}a^{\prime})=\operatorname{tp}(\bar{b}^{\prime}a^{\prime\prime})$. Let $g\in G$ be such that $g\cdot a=a^{\prime}$ and let $E^{\prime}=g\cdot E$. In particular, $E^{\prime}\in\mathcal{F}_{a^{\prime}}$. Let $B_{n}^{\prime}\subseteq B^{\prime}$ be such that there is $h_{n}\in G$ satisfying $h_{n}\cdot B_{n}=B_{n}^{\prime}$ and $h_{n}\cdot a=a^{\prime}$. This implies that $h_{n}\cdot E=E^{\prime}$. Let $E_{n}^{\prime}=h_{n}\cdot E_{n}$, so that we have $\mu(E_{n}^{\prime}\triangle E^{\prime})<1/n$ and $E_{n}^{\prime}\in\mathcal{F}_{B_{n}^{\prime}}$. Let $g_{n}\in G_{(B_{n})}$ be such that $g_{n}\cdot a^{\prime}=a^{\prime\prime}$, so that $g_{n}\cdot E^{\prime}\in\mathcal{F}_{a^{\prime\prime}}$. We have:
@@
Suppose now that missing (ii) holds and let $d$ be such that for every $h\notin G_{a}$, every element of $G$ can be written as a word of length at most $d$ in $h,h^{-1}$, and elements of $G_{a}$. Let $\epsilon>0$ and let $n>(4d^{2})/\epsilon^{2}$. By the no algebraicity assumption, there is $h\in G_{(B_{n})}$, $h\notin G_{a}$. Let $E^{\prime}=h\cdot E$, so that $\mu(E\triangle E^{\prime})<2/n$. Let $\xi={\mathbf{1}}_{E}$ and note that
@@
Suppose that $G\leq\operatorname{Sym}(M)$ is a primitive, oligomorphic permutation group with no algebraicity. Let $G\curvearrowright[0,1]^{M}$ be the shift action. Then any $G$-invariant, ergodic measure on $[0,1]^{M}$ is of the form $\lambda^{\otimes M}$, where $\lambda$ is a probability measure on $[0,1]$.
@@
If the permutation group $G\leq\operatorname{Sym}(M)$ is oligomorphic, it satisfies both conditions missing (i) and missing (ii) of Section 4. Indeed, missing (i) follows from the no algebraicity assumption and the compactness theorem for first-order logic (by considering $M$ as an $\aleph_{0}$-categorical structure) and missing (ii) follows from primitivity and the fact that each $G_{a}$ has only finitely many double cosets in $G$. ∎

### v2

@@ 3.2. Induced actions
The action does not depend on the choice of the section $s$ (up to conjugacy). Note that if $\sigma$ preserves the measure $\mu$, then $\operatorname{Ind}_{V}^{G}(\sigma)$ preserves the measure $\xi\otimes\mu$. Also, if $\sigma$ is ergodic, then so is $\operatorname{Ind}_{V}^{G}(\sigma)$.
@@ 3.3. A classification theorem for non-singular actions
First, we show that there is a non-null $G$-invariant set $E\in\mathrm{MALG}(\mu)$ such that the action of $G$ on $E$ is isomorphic to $\operatorname{Ind}_{V}^{G}(\alpha)$ as above. Let $\mathcal{H}=L^{2}(\Omega,\mu)$ and consider the Koopman representation $\kappa$ of $G$ on $\mathcal{H}$ given by (3.1). Let $(\mathcal{A},\leq)$ and the subspaces $\mathcal{H}_{a}$ for $a\in\mathcal{A}$ be defined as in Subsection 2.2. By Subsection 2.2, $\bigcup_{a\in\mathcal{A}}\mathcal{H}_{a}$ is dense in $\mathcal{H}$, so there exists $a$ with $\mathcal{H}_{a}\neq 0$. By Subsection 2.1, there is a <<-minimal element $a\in\mathcal{A}$ with $\mathcal{H}_{a}\neq 0$.
@@
Denote the probability measure-preserving action of $G_{a}$ on $(E,\nu_{a})$ by $\alpha$. We will see that the actions $\operatorname{Ind}_{G_{a}}^{G}(\alpha)$ and $G\curvearrowright\bigsqcup_{t\in T}t\cdot E$ are isomorphic. In order to make the isomorphism more transparent, we choose the section $s\colon G/G_{a}\to G$ appearing in the definition of the cocycle () 3.2 to have image $T$, i.e,
@@
To conclude with the proof of the theorem, we apply Zorn’s lemma. Let $\mathcal{D}$ be the collection of all elements of $\mathrm{MALG}(\mu)$ which are $G$-invariant and such that the restriction of the action of $G$ to them is of the form (3.3), and order it by inclusion. If $\mathcal{C}\subseteq\mathcal{D}$ is a chain, we can always find a cofinal set $\mathcal{C}^{\prime}\subseteq\mathcal{C}$, which is countable, because $(\mathcal{C},\subseteq)$ is isomorphic to $(\{\mu(C):C\in\mathcal{C}\},\leq)\subseteq({\mathbf{R}},\leq)$. It is clear that $\bigcup\mathcal{C}^{\prime}$ is an upper bound for $\mathcal{C}$. Let $D$ be a maximal element of $\mathcal{D}$. If $D\neq\Omega$, we can apply what we already proved to the action of $G$ on $\Omega\setminus D$ and contradict its maximality. ∎
@@ 4. A generalization of de Finetti’s theorem
Let $E\in\mathcal{J}$. It follows from () 4.1 that there exists a finite set $A\subseteq M$ and a $\mathcal{F}_{A}$-measurable event $E^{\prime}$ with $\mu(E\triangle E^{\prime})<\epsilon$. If $B\subseteq M$ is any finite set, by Section 4, there is $g\in G$ with $g\cdot A\cap B=\emptyset$. By invariance of $E$, this implies that for any $B$, there is an $\mathcal{F}_{M\setminus B}$-measurable event $E^{\prime\prime}$ (namely, $E^{\prime\prime}=g\cdot E^{\prime}$) with $\mu(E\triangle E^{\prime\prime})<\epsilon$. Taking limits as $B$ exhausts $M$ and $\epsilon\to 0$, we obtain that $E\in\mathcal{T}$. ∎
@@
Let $\epsilon>0$. By reverse martingale convergence [Durrett2005]*Theorem 5.6.3, we have $\operatorname{\mathbf{E}}_{\mathcal{G}_{n}}\xi\to\operatorname{\mathbf{E}}_{\mathcal{T}}\xi$ in $L^{2}$; let $n$ be such that $\lVert\operatorname{\mathbf{E}}_{\mathcal{G}_{n}}\xi\rVert-\lVert\operatorname{\mathbf{E}}_{\mathcal{T}}\xi\rVert<\epsilon$ (here $\lVert\cdot\rVert$ is the $L^{2}$-norm). We have:
@@
For every $g\in G_{a_{i}}$, we have that $g\cdot\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}=\operatorname{\mathbf{E}}_{g\cdot\mathcal{T}}g\cdot\xi_{i}=\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}$, so $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{i}\in\mathcal{F}_{a_{i}}$. In particular, $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}$ is $\mathcal{G}$-measurable, so also $\mathcal{T}\cap\mathcal{G}$-measurable. As $\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}\operatorname{\mathbf{E}}_{\mathcal{T}}=\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}$, this implies that $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}=\operatorname{\mathbf{E}}_{\mathcal{T}\cap\mathcal{G}}\xi_{1}\cdots\xi_{n}$. Linear combinations of variables of the form $\xi_{1}\cdots\xi_{n}$ generate a dense subspace of $L^{2}(\mathcal{G})$, so $\operatorname{\mathbf{E}}_{\mathcal{T}}\xi_{1}\cdots\xi_{n}$ generate a dense subspace of $L^{2}(\mathcal{T}\cap\mathcal{G})$. However, by () 4.2, each of them belongs to $\bigvee_{a}\mathcal{T}\cap\mathcal{F}_{a}$ and we are done. ∎
@@
Suppose now that (i) holds and let the tuple $\bar{b}$ enumerate the set $\bigcup_{n}B_{n}$. By uniform non-algebraicity, there exist $\bar{b}^{\prime}$ and $a^{\prime}\neq a^{\prime\prime}$ with $\operatorname{tp}(\bar{b}a)=\operatorname{tp}(\bar{b}^{\prime}a^{\prime})=\operatorname{tp}(\bar{b}^{\prime}a^{\prime\prime})$. Let $g\in G$ be such that $g\cdot a=a^{\prime}$ and let $E^{\prime}=g\cdot E$. In particular, $E^{\prime}\in\mathcal{F}_{a^{\prime}}$. Let $B_{n}^{\prime}\subseteq B^{\prime}$ be such that there is $h_{n}\in G$ satisfying $h_{n}\cdot B_{n}=B_{n}^{\prime}$ and $h_{n}\cdot a=a^{\prime}$. This implies that $h_{n}\cdot E=E^{\prime}$. Let $E_{n}^{\prime}=h_{n}\cdot E_{n}$, so that we have $\mu(E_{n}^{\prime}\triangle E^{\prime})<1/n$ and $E_{n}^{\prime}\in\mathcal{F}_{B_{n}^{\prime}}$. Let $g_{n}\in G_{(B_{n})}$ be such that $g_{n}\cdot a^{\prime}=a^{\prime\prime}$, so that $g_{n}\cdot E^{\prime}\in\mathcal{F}_{a^{\prime\prime}}$. We have:
@@
Suppose now that (ii) holds and let $d$ be such that for every $h\notin G_{a}$, every element of $G$ can be written as a word of length at most $d$ in $h,h^{-1}$, and elements of $G_{a}$. Let $\epsilon>0$ and let $n>(4d^{2})/\epsilon^{2}$. By the no algebraicity assumption, there is $h\in G_{(B_{n})}$, $h\notin G_{a}$. Let $E^{\prime}=h\cdot E$, so that $\mu(E\triangle E^{\prime})<2/n$. Let $\xi={\mathbf{1}}_{E}$ and note that
@@
Suppose that $G\leq\operatorname{Sym}(M)$ is a primitive, oligomorphic permutation group with no algebraicity. Let $G\curvearrowright[0,1]^{M}$ be the shift action. Then any $G$-invariant, ergodic measure on $0,1]^{M}$ is of the form $\lambda^{\otimes M}$, where $\lambda$ is a probability measure on $[0,1]$.
@@
If the permutation group $G\leq\operatorname{Sym}(M)$ is oligomorphic, it satisfies both conditions (i) and (ii) of Section 4. Indeed, (i) follows from the no algebraicity assumption and the compactness theorem for first-order logic (by considering $M$ as an $\aleph_{0}$-categorical structure) and (ii) follows from primitivity and the fact that each $G_{a}$ has only finitely many double cosets in $G$. ∎
@@
We end the paper with several examples of non-oligomorphic groups where Section 4 applies. All of the examples are constructed as appropriate Fraïssé limits.
Example 4.13.
Let ${\mathbf{U}}_{\mathbf{Z}}$ be the integer Urysohn space, i.e., the Fraïssé limit of all finite metric spaces with integer-valued metric. It is well-known that this is an amalgamation class and it is not difficult to check that the action $\operatorname{Iso}({\mathbf{U}}_{\mathbf{Z}})\curvearrowright{\mathbf{U}}_{\mathbf{Z}}$ satisfies condition (i) and the other hypotheses of Section 4. However, it does not satisfy condition (ii) (and it is not oligomorphic). Indeed, let $a\in{\mathbf{U}}_{\mathbf{Z}}$ be any point and let $g\notin G_{a}$. Then any element of $G$ which can be written as a word of bounded length in elements of $G_{a}$ and $g,g^{-1}$ can move the point $a$ only a bounded distance from itself. Section 4 for ${\mathbf{U}}_{\mathbf{Z}}$ is not new—it was already proved in [Barritault2024pp]—but this example is a good illustration for the type of situations where condition (ii) fails.
Example 4.14.
Let $\mathcal{L}=\{R_{i}:i\in{\mathbf{N}}\}$ be a signature with infinitely many binary relations. Let $\mathcal{F}_{1}$ be the Fraïssé class of all finite $\mathcal{L}$-structures such that all relations are symmetric and for every pair of points, exactly one relation holds. Let $M_{1}$ be its Fraïssé limit. Then the action $\operatorname{Aut}(M_{1})\curvearrowright M_{1}$ is not oligomorphic (because it has infinitely many orbits on pairs) and it satisfies both conditions (i) and (ii) of Section 4.
Next consider the Fraïssé class $\mathcal{F}_{2}$ in the same signature where, in addition, one forbids monochromatic triangles, i.e., configurations $\{R_{i}(a,b),R_{i}(b,c),R_{i}(c,a)\}$ for all $i\in{\mathbf{N}}$, and let $M_{2}$ denote the Fraïssé limit. Then condition (ii) holds but (i) fails. To see this, let $a\in M_{2}$ and let $\bar{b}$ be an infinite tuple such that $R_{i}(a,b_{i})$ for all $i$. Then it is impossible to have two distinct points $a^{\prime},a^{\prime\prime}$ which have the type of $a$ over a copy $\bar{b}^{\prime}$ of $\bar{b}$: $R_{i}(a^{\prime},a^{\prime\prime})$ must hold for some $i$ and this will create a monochromatic triangle $\{a^{\prime},a^{\prime\prime},b_{i}^{\prime}\}$.